OpenStudy (anonymous):

(8.3 x 10^7)(5.9 x 10^8) in scientific notation

OpenStudy (anonymous):

(8.3 x 10^7)(5.9 x 10^8) =(8.3 x 5.9)(10^15) =4.90 x 10^16

OpenStudy (anonymous):

8.3 * 5.9 = 48.97 10^7 * 10^8 = 10 ^ 15 divide 48.97 by 10 and * 10^15 by 10 = 4.897* 10^16